Shanti drove 78 miles in 1 1/2 hours or (1 and 1 half). At the same rate, how many miles could she drive in 2.5 hours?
Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
r=d/t=78/(1.5)=52 mph
d=rt=(52m/h)(2.5h)=130 mi
Hope this helps---ptaylor